# Can MLLMs Guide Me Home? A Benchmark Study on Fine-Grained Visual Reasoning from Transit Maps

### v2

@@ 4.1 Correctness Evaluation
We evaluate the correctness of the answer using Algorithm 1 in Appendix B. Specifically, the evaluation checks the correctness of the overall departure and arrival stops (stop1 and stop2), verifies if the route name of each segment exists in the Map Metadata, ensures the departure and arrival stops are valid for each segment, and confirms that transfer stops between consecutive segments are consistent. An answer is considered correct only if all the above checks are satisfied. Additionally, we apply the same correctness evaluation algorithm to the answers of short and long questions.
@@ 4.2 Quality Evaluation
For short questions, the map score solely focuses on route-level and endpoint consistency, excluding all long-question-specific parts. We compute the score by comparing segment pairs in the answer and reference route. Specifically, correctly matching stop1 and stop2 contributes one point, matching the route name adds two points, and matching the departure and arrival stops within each route segment provides one point each. The final score is capped at $10$, and an additional bonus is awarded if the answer is judged correct based on the correctness evaluation procedure described in Section 4.1. This design ensures that a correct answer always receives a higher score than any incorrect one.
For long questions, the evaluation incorporates additional scoring components tailored to the two question templates introduced in Section 3.1.2. These components are designed to capture the increased reasoning depth required in long-form responses. As with short questions, a bonus score is also added for correct answers. The two additional scoring components are detailed below.
@@
Specific Via Stop Evaluation. For long questions that require explicit enumeration of intermediate stops, we compute via_stop_score using a combination of two factors: the number of correctly matched via stops, and the intersection-over-union (IoU) between via stop sets of the answer and reference route. The final score for this component is obtained by averaging the IoU score (scaled to $10$) and the exact match count (capped at $10$), and then clipped to a maximum of $10$.
5 Experiments
@@
Qwen2.5-VL-32B-Instruct [2] Base $16.49\%$ $36$ $15.71\%$ $112$ $3.88$ / $6.84$
@@
Kimi-VL-A3B-Thinking [16] Reasoning $5.47\%$ $754$ $5.47\%$ $1,287$ $2.44$ / $3.17$
@@ 5.1 Experimental Setups
We conduct extensive benchmark experiments on ReasonMap using $15$ popular MLLMs under different inference settings, analyzing their performance and comparing results. Several interesting insights emerge from this comparison. The detailed experimental settings are described below.
Evaluated Models. We evaluate a diverse set of MLLMs categorized into two groups based on whether they are reasoning-oriented models with a long-thinking process. Reasoning models include: Skywork-R1V-38B [17, 18], QvQ-72B-Preview [20], Kimi-VL-A3B-Thinking/Instruct [16], OpenAI o3 [48], Gemini-2.5-Flash [50], Doubao-1-5-thinking-vision-pro-250428 (Doubao-428), and Doubao-1.5-Thinking-Pro-M-250415 (Doubao-415) [19]. Base models include: Qwen2.5-VL series (3B, 32B, 72B) [2], InternVL3 series (38B, 78B) [3], OpenAI 4o [49], and Doubao-1.5-Vision-Pro-32k-250115 (Doubao-115) [19]. Additionally, the Doubao 1.5 Pro series has an activated parameter size of $20$B.
Inference Settings. For open-source models, we set the max output token limit to $2,048$, while keeping other parameters consistent with the official HuggingFace configurations. All open-source models are deployed using PyTorch and the HuggingFace Transformers library on $8$ NVIDIA A100 GPUs. For closed-source models, we access their official APIs for evaluation and follow the default settings provided by each model’s official documentation. We further discuss the diverse image processing strategies when handling high-resolution visual inputs in Appendix D.
Difficulty-Aware Weighting. To better reflect the varying complexity of different samples, we adopt a difficulty-aware weighting strategy based on the combination of question difficulty and map difficulty. Specifically, each difficulty pair is assigned a predefined weight, with harder combinations receiving higher values. The complete weight matrix is provided in Appendix B.2. Both accuracy and map score are evaluated using this weighting scheme, ensuring that models are more strongly rewarded for correctly solving more challenging examples.
@@ 5.2 Experimental Results
5.2.1 Performance of MLLMs with Full Input
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure 3: Accuracy across difficulty combinations for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each difficulty combination is denoted by a pair (e.g., easy-hard), where the first term indicates question difficulty and the second term represents map difficulty. The pair (hard-middle) contains only one sample, leading to an accuracy of 100%. We summarize the number of evaluation samples in each difficulty bucket: $55$ samples for easy-easy, $46$ for easy-middle, $28$ for middle-easy, $7$ for hard-easy, $23$ for middle-middle, $80$ for easy-hard, $1$ for hard-middle, $57$ for middle-hard, and $15$ for hard-hard.
The principal results are summarized in Table 2. Notably, we observe a counterintuitive phenomenon: among open-source models, reasoning models consistently underperform their base counterparts, whereas the opposite holds in the closed-source setting [^3] Although the comparison across closed-source models may not be fair due to lack of transparency in details, the reasoning variants exhibit stronger performance in this category.. Prior work suggests that RL may improve sample efficiency without introducing fundamentally new reasoning abilities [59, 60, 61], while RL-trained models tend to bias their output distributions toward high-reward trajectories, which helps produce more correct responses but may simultaneously constrain the model’s exploration capacity and reduce its ability to leverage broader foundational knowledge. In addition, recent studies indicate that multimodal models may sometimes rely on inner knowledge priors instead of truly attending to visual inputs [62, 63, 64, 30]. This tendency is further supported by the results in Section 5.2.2, where open-source models still maintain part of their performance even without visual input, indicating limited visual grounding. In contrast, closed-source reasoning models outperform their base variants. One possible explanation lies in the broader knowledge coverage and better visual integration observed in these models [19, 48, 50].
We further analyze the effect of model size by examining performance within the same architecture series. Qwen2.5-VL and InternVL series show a consistent trend: larger models achieve better accuracy with fewer tokens, suggesting that the scaling law [65] continues to hold even in fine-grained visual reasoning tasks. Figure 3 presents accuracy distributions across different combinations of question and map difficulty. As expected, performance degrades as task complexity increases. Additionally, Figure 4 illustrates accuracy variation across cities. We observe a negative correlation between map difficulty and accuracy. Moreover, model performance varies notably even among cities with comparable map difficulty levels. This disparity can be partially attributed to factors such as city prominence and the language used for stop names, both of which are closely tied to the model’s pretrained knowledge. For instance, OpenAI o3 performs significantly better on complex cities like Singapore compared to Hangzhou, likely due to Singapore’s higher international visibility and the use of English stop names, whereas Hangzhou is less prominent and its stop names are Chinese.
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure 4: Accuracy across different cities for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each city is marked with the corresponding map difficulty and the country flag. Each city in the test set provides a specific number of samples per model: $32$ samples for Auckland, $34$ for Los Angeles, $7$ for Miami, $35$ for Lisboa, $18$ for Geneva, $40$ for Beijing, $39$ for Hangzhou, $17$ for Budapest, $39$ for Singapore, $40$ for Rome, and $11$ for Toronto.
Table 3: Evaluations of various MLLMs on ReasonMap without visual inputs. $S.$ represents results for short questions, while $L.$ denotes results for long questions. The map score is capped at 20 for short questions, while for long questions, the maximum score is 40. Bold indicates the best results among open-source and closed-source models, respectively, while underline represents the second best. Green highlights improved results compared to the full input setting (Table 2), while red indicates performance drops.
@@ 5.2.1 Performance of MLLMs with Full Input
[…]
@@
[…]
@@ 5.2.2 Performance of MLLMs without Visual Input
To further investigate the reliance of MLLMs on visual input, we selected representative open-source and closed-source models for additional experiments, where the visual input was masked. The results are reported in Table 3. We observe that while most models can leverage internal knowledge to answer certain questions, their performance generally declines to varying degrees when visual input is removed, with the decline being more pronounced among closed-source models. Model performance is positively correlated with the performance drop after masking visual inputs, indicating effective use of visual information. In contrast, models like Qwen2.5-VL-3B-I show minimal or even improved performance, suggesting a reliance on inner knowledge rather than genuine visual reasoning.
@@ 5.3 Error Analysis
Figure 5: Error case analysis of various MLLMs using ReasonMap. For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and categorize them accordingly.
Figure 5 presents representative failure cases from ReasonMap, revealing several recurring error types. A common issue is visual confusion, where the model misidentifies the transit line due to similar colors or adjacent layouts, for instance, mistaking Line $9$ for Line $16$ (OpenAI o3, left column; Doubao-428, right column). Another frequent problem is format errors, where responses deviate from the required structure, making them unprocessable despite containing correct route information (Doubao-115 and QvQ-72B-Preview, left column). We also observe instances of hallucination [58], where the model repeats the correct answer (Kimi-VL-A3B-Thinking, middle column) or generates information that is not present in the input, such as mentioning image generation, as seen in Skywork-R1V-38B (right column). Refusal cases are also present, where models explicitly decline to answer (Skywork-R1V-38B, middle and right column). Notably, these errors may occasionally co-occur within a single response (Skywork-R1V-38B, right column). These behaviors highlight the limitations in visual grounding and response robustness, especially when handling fine-grained visual details (see more case analyses in Appendix C).
@@ 6 Conclusion
In this work, we introduce ReasonMap, a benchmark designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. Through a semi-automated and scalable data building pipeline, we curate a diverse set of human-verified question-answer pairs across $30$ cities. Our two-level evaluation framework enables a nuanced assessment of both correctness and quality. Experimental results on $15$ popular MLLMs reveal key insights into model behavior, highlighting performance gaps between base and reasoning models, as well as the crucial role of visual input. These findings underscore the need for more rigorous evaluation and training approaches to advance visual reasoning in multimodal systems.

### v3

@@ 4.1 Correctness Evaluation
We evaluate the correctness of the answer using Algorithm 1 in Appendix B. Specifically, the evaluation checks the correctness of the overall departure and arrival stops (stop1 and stop2), verifies if the route name of each segment exists in the Metro Data, ensures the departure and arrival stops are valid for each segment, and confirms that transfer stops between consecutive segments are consistent. An answer is considered correct only if all the above checks are satisfied. Additionally, we apply the same correctness evaluation algorithm to the answers of short and long questions.
@@ 4.2 Quality Evaluation
For short questions, the map score solely focuses on route-level and endpoint consistency, excluding all long-question-specific parts. We compute the score by comparing segment pairs in the answer and reference route. Specifically, correctly matching stop1 and stop2 contributes one point, matching the route name adds two points, and matching the departure and arrival stops within each route segment provides one point each. The final score is capped at $10$, and an additional bonus is awarded if the answer is judged correct based on the correctness evaluation procedure described in Section 4.1. This design ensures that a correct answer always receives a higher score than any incorrect one. For long questions, the evaluation incorporates additional scoring components tailored to the two question templates introduced in Section 3.1.2. These components are designed to capture the increased reasoning depth required in long-form responses. As with short questions, a bonus score is also added for correct answers. The two additional scoring components are detailed below.
@@
Specific Via Stop Evaluation. For long questions that require explicit enumeration of intermediate stops, we compute via_stop_score using a combination of two factors: the number of correctly matched via stops, and the intersection-over-union (IoU) between via stop sets of the answer and reference route. The final score for this component is obtained by averaging the IoU score (scaled to $10$) and the exact match count (capped at $10$), and then clipped to $10$.
4.3 Training Baseline
Building upon the evaluation framework established above, we introduce a strong training baseline. This baseline serves not only to validate the efficacy of our proposed metrics but also to establish a robust performance benchmark for the ReasonMap. The metrics defined in our framework, such as route correctness and map score (see in Section 4.1 and 4.2), are rule-based and non-differentiable. Consequently, standard supervised fine-tuning (SFT) is insufficient as it cannot directly optimize for these desired outcomes. Therefore, we leverage reinforcement learning (RL) to explicitly align the model outputs with our evaluation criteria.
Specifically, we fine-tune Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct [3] on the ReasonMap training set using RL via the GRPO procedure [57] (see detailed GRPO optimization process in Appendix B.4). The reward is designed in accordance with our evaluation framework, consisting of two components: (1) accuracy reward, which provides a binary signal following the correctness evaluation; and (2) format reward, which encourages parsable outputs by rewarding correct formats and penalizing violations. The resulting RL-trained models are evaluated across unseen cities, enabling a cross-city validation of generalization ability (see results in Section 5.2.3).
@@ 4.3 Training Baseline
Qwen2.5-VL-7B-Instruct [3] Base $13.28\%$ $26$ $7.12\%$ $57$ $4.01$ / $5.74$
Qwen2.5-VL-32B-Instruct [3] Base $16.49\%$ $36$ $5.71\% 11$2112 $3.88$ / $6.84$
@@
Kimi-VL-A3B-Thinking [64] Reasoning $\% 754754$ $5.47\%$ $1,287$ $2.44$ / $3.17$
@@ 5.1 Experimental Setups
We conduct extensive benchmark experiments on ReasonMap using $16$ popular MLLMs under different inference settings, analyzing their performance and comparing results. Several interesting insights emerge from this comparison. The detailed experimental settings are as follows.
Evaluated Models. We evaluate a diverse set of MLLMs categorized into two groups based on whether they are reasoning-oriented models with a long-thinking process. Reasoning-oriented models include: Skywork-R1V-38B [74, 49], QvQ-72B-Preview [51], Kimi-VL-A3B-Thinking/Instruct [64], OpenAI o3 [48], Gemini-2.5-Flash [19], Doubao-1-5-thinking-vision-pro-250428 (Doubao-428), and Doubao-1.5-Thinking-Pro-M-250415 (Doubao-415) [6]. Base models include: Qwen2.5-VL series (3B, 7B, 32B, 72B) [3], InternVL3 series (38B, 78B) [102], OpenAI 4o [46], and Doubao-1.5-Vision-Pro-32k-250115 (Doubao-115) [6]. Additionally, the Doubao 1.5 Pro series has an activated parameter size of 2020B.
Inference Settings. For open-source models, we set the max output token limit to $2,048$, which is sufficient to cover complete generations, even for reasoning-oriented models, while keeping other parameters consistent with the official HuggingFace configurations. All open-source models are deployed using PyTorch and the HuggingFace Transformers library. For closed-source models, we access their official APIs for evaluation and follow the default settings provided by each model’s official documentation. We further discuss the diverse image processing strategies when handling high-resolution visual inputs in Appendix B.2.
Training Settings. For GRPO-based RL training, we use AdamW with an initial learning rate of $1.0\times 10^{-6}$ and a KL divergence coefficient of $1.0\times 10^{-3}$. We sample $8$ responses every query and set the global batch size to $16$. The training data is the training set from ReasonMap.
Difficulty-Aware Weighting. To emphasize models that are capable of addressing more challenging cases, we incorporate a difficulty-aware weighting scheme when computing the metrics. Samples with higher map or question difficulty are assigned larger weights, allowing the evaluation to more faithfully reflect a model’s robustness. The complete weighting details are provided in Appendix B.3.
@@ 5.2 Experimental Results
Table 3: Evaluations of MLLMs on ReasonMap w/o visual inputs. S.{S.} denotes results for short questions and L.{L.} denotes results for long questions. The map score is capped at $20$ for short questions, while for long questions, the maximum score is $40$. Bold indicates the best results among open-source and closed-source models, respectively, while underline represents the second best. Green highlights improved results compared to the full input setting (Table 2), while red indicates performance drops.
@@
Qwen2.5-VL-3B-Instruct [3] Base $9.38\%$ $\uparrow\!$ $0.7\%$ $47$ $9.72\%$ $\uparrow\!$ $1.73\%$ $147$ $2.93$ $\uparrow\!$ $0.18$ / $4.51$ $\uparrow\!$ $0.81$
Qwen2.5-VL-72B-Instruct [3] Base $\mathbf{16.41\%}$ $\downarrow\!$ $10.24\%$ $28$ $\mathbf{15.71\%}$ $\downarrow\!$ $8.51\%$ $108$ $\mathbf{4.03}$ $\downarrow\!$ $1.06$ / $\mathbf{6.49}$ $\downarrow\!$ $2.31$
Kimi-VL-A3B-Instruct [64] Base $\underline{11.81\%}$ $\downarrow\!$ $0.95\%$ $41$ $\underline{9.81\%}$ $\downarrow\!$ $2.52\%$ $49$ $\underline{3.37}$ $\uparrow\!$ $0.07$ / $\underline{5.32}$ $\downarrow\!$ $0.05$
Kimi-VL-A3B-Thinking [64] Reasoning $4.17\%$ $\downarrow\!$ $1.30\%$ $1,039$ $2.08\%$ $\downarrow\!$ $3.39\%$ $1,755$ $2.06$ $\downarrow\!$ $0.38$ / $1.64$ $\downarrow\!$ $1.53$
@@
Doubao-115 [6] Base $\underline{13.72\%}$ $\downarrow\!$ $20.48\%$ $34$ $\underline{13.98\%}$ $\downarrow\!$ $24.04\%$ $99$ $\underline{3.50}$ $\downarrow\!$ $1.75$ / $\underline{6.48}$ $\downarrow\!$ $5.48$
Doubao-415 [6] Reasoning $\mathbf{21.53\%}$ $\downarrow\!$ $21.61\%$ $352$ $\mathbf{17.19\%}$ $\downarrow\!$ $28.90\%$ $1,047$ $\mathbf{4.85}$ $\downarrow\!$ $2.48$ / $\mathbf{7.68}$ $\downarrow\!$ $6.99$
5.2.1 Performance of MLLMs with Full Input
The principal results are summarized in Table 2 (see more fine-grained metrics in Table A3 of the Appendix). We first analyze the effect of model size by examining performance within the same architecture series. Qwen2.5-VL and InternVL series show a consistent trend: larger models achieve better accuracy with fewer tokens, suggesting that the scaling law [33] continues to hold even in fine-grained visual reasoning tasks. Figure A1 in Appendix presents accuracy distributions across different combinations of question and map difficulty. As expected, performance degrades as task complexity increases. Additionally, Figure A2 in Appendix illustrates accuracy variation across cities. We observe a negative correlation between map difficulty and accuracy. Moreover, model performance varies notably even among cities with comparable map difficulty levels. This disparity can be partially attributed to factors such as city prominence and the language used for stop names (see the ablation study on language in Appendix C.3), both of which are closely tied to the model’s pretrained knowledge.
Notably, we observe a counterintuitive phenomenon: among open-source models, reasoning models consistently underperform their base counterparts, whereas the opposite holds in the closed-source models [^1] Although the comparison across closed-source models may not be fair due to lack of transparency in details, the reasoning variants exhibit consistently stronger performance in this category.. Recent works argue that RL improves sample efficiency rather than fostering new reasoning abilities [89, 73, 95], as it biases models toward high-reward responses, limits exploration, and, because most open-source reasoning models are fine-tuned mainly on math, physics, and logic tasks [57, 49], further reinforces domain-specific reasoning behaviors. However, these do not fully explain why closed-source reasoning models outperform their base variants. To further investigate, we conduct a case analysis. Among open-source models, we observe that reasoning variants often engage in repeated trial-and-verification during their thinking process. This iterative reasoning frequently introduces visual confusion; for instance, a model may initially identify the correct route but later override it with an incorrect one during self-reflection, failing to recover the earlier correct reasoning (see the case by Kimi-VL-A3B-Thinking). In contrast, closed-source models exhibit stronger visual grounding: even when visual confusion occurs, they tend to self-correct within the reasoning chain, ultimately producing the correct answer as illustrated in Figure A3 of the Appendix. Overall, this suggests that the better performance of closed-source reasoning models is primarily driven by their stronger visual grounding capability, which enables effective error correction throughout the reasoning process.
Wrong Case by Kimi-VL-A3B-Thinking (green denotes correct & red represents wrong) <think><think>Is Augustins on Line 18? Yes, it is on Line 18. … Wait, maybe I got the location wrong … Let me confirm, steps should be like: Augustins (Line 14) - … </think></think>Route: Line 14
@@ 5.2.2 Performance of MLLMs without Visual Input
Recent studies suggest that multimodal models may rely on internal knowledge priors rather than truly attending to visual inputs [31, 24, 20, 92]. To examine this phenomenon, we further evaluate representative open- and closed-source models under the visual-masking setting. The results are reported in Table 3. We observe that while most models can leverage prior knowledge to answer questions, their performance generally declines to varying degrees when visual input is removed, with the decline being more pronounced among closed-source models. Model performance is positively correlated with the performance drop after masking visual inputs, indicating effective use of visual information. In contrast, models like Qwen2.5-VL-3B-I show minimal or even improved performance, suggesting a reliance on prior knowledge rather than real visual reasoning. We further conduct non-vision experiments by replacing maps with their symbolic representations in Appendix C.4.
5.2.3 Results of Training Baseline
Table 4: Evaluations of RL-trained models on ReasonMap. $S.$ represents results for short questions, while $L.$ denotes results for long questions. The map score is capped at $20$ for short questions, while for long questions, the maximum score is $40$.
Qwen2.5-VL-3B-Instruct [3] Base $8.68\%$ $42$ $7.99\%$ $151$ $2.75$ / $3.70$
+ Reinforcement Fine-tuning Base $11.46\%$ $\uparrow\!$ $2.78\%$ $25$ $10.50\%$ $\uparrow\!$ $2.51\%$ $93$ $3.81$ $\uparrow\!$ $1.06$ / $6.09$ $\uparrow\!$ $2.39$
Qwen2.5-VL-7B-Instruct [3] Base $13.28\%$ $26$ $7.12\%$ $57$ $4.01$ / $5.74$
+ Reinforcement Fine-tuning Base $26.22\%$ $\uparrow\!$ $12.94\%$ $25$ $26.04\%$ $\uparrow\!$ $18.92\%$ $34$ $5.52$ $\uparrow\!$ $1.51$ / $9.52$ $\uparrow\!$ $3.78$
As shown in Table 4, the training baseline with the GRPO-based RL scheme consistently improves overall accuracy and map reasoning quality across model scales, while also reducing token usage. These results demonstrate that incorporating the reward signals can effectively enhance reasoning efficiency and improve answer quality.
@@ 5.3 Error Analysis
Figure 3: Error case analyses of various MLLMs using ReasonMap. For reasoning models, the reasoning process is explicitly marked with <think> and </think> tags. We highlight error contents in the answers with red and categorize them accordingly.
Figure 3 presents representative failure cases from ReasonMap, revealing several recurring error types. A common issue is visual confusion, where the model misidentifies the transit line due to similar colors or adjacent layouts, for instance, mistaking Line $9$ for Line $16$ (OpenAI o3, left column; Doubao-428, right column). Another frequent problem is format errors, where responses deviate from the required structure, making them unprocessable despite containing correct route information (Doubao-115 and QvQ-72B-Preview, left column). We also observe instances of hallucination [4], where the model repeats the correct answer (Kimi-VL-A3B-Thinking, middle column) or generates information that is not present in the input, such as mentioning image generation, as seen in Skywork-R1V-38B (right column). Refusal cases are also present, where models explicitly decline to answer (Skywork-R1V-38B, middle and right column). Notably, these errors may occasionally co-occur within a single response (Skywork-R1V-38B, right column). Furthermore, we conduct a systematic analysis of failure causes from a model capability perspective (e.g., optical character recognition (OCR), visual grounding, and spatial reasoning) in Appendix C.5. The above error types highlight the limitations in visual grounding and response robustness, especially when handling fine-grained visual details (see more cases in Appendix D).
@@ 6 Conclusion
In this work, we introduce ReasonMap, a new benchmark specifically designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. Through a semi-automated and scalable data building pipeline, we curate a diverse set of human-verified question-answer pairs across $30$ cities from $13$ countries. Our two-level evaluation framework enables a nuanced assessment of both correctness and quality. Comprehensive experiments on $16$ widely used MLLMs reveal key insights into model behavior, highlighting performance gaps between base and reasoning models, as well as the crucial role of visual input. To further strengthen the benchmark and facilitate future research, we provide a GRPO-based RL baseline. Error analyses also reveal recurring failure patterns, highlighting visual weaknesses of current MLLMs. Collectively, these findings underscore the need for more rigorous evaluation and training approaches to advance visual reasoning.
Acknowledgement
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and Hangzhou Municipal Talent Program.
Supplementary Material